6. Hydrex Mechanical Products is planning to set aside S160,000 now whenever it becomes necessary. If the replacement is expected to take place in 3.5 years, how much will the company have in its investment set-aside account? Assume a rate of return of 4% per quarter? (ANSWER: $277,072)
Answer #1

Here, P = 160,000,       i = 4%.      n = 3.5 * 4 = 14 quarters

Future Value = P(F/P, i, n)

                     = 160,000(F/P, 4%, 14)

                     = 160,000(1.7317)

                     = $277,072

Thus, the company will have $222,072 in its account